OpenStudy (anonymous):

why does L'hospital's rule fail for evaluating this limit: x tends to 0, [e^(1/x) - 1]/[e^(1/x) + 1]. The RHL is not equal to LHL clearly limit doesnt exist however LH rule yields the value of RHL(1). why??

OpenStudy (anonymous):

I found this at matharticles.com on L'Hopitals rule that might shed some light on your problem. http://www.maa.org/pubs/calc_articles/ma041.pdf Good luck, TTD

OpenStudy (anonymous):

the function isn't continues at x=0 and even isn't defined at 0 so you can't calculate its derivative and use that in L'Hopital's rule at x=0 (its graph is could be shaped from tanhx, with 1/x=y substituion)

OpenStudy (anonymous):

@arnp5 LHospitals do apply since....U will get inderterminant of the form \[\infty/\infty\]

OpenStudy (anonymous):

and the answer is 1